# Semantic Complete Scene Forecasting from a 4D Dynamic Point Cloud Sequence

### 13 Dec 2023 v1

@@ 1 Introduction
The SCSF problem is very challenging since it requires both future forecasting and semantic scene completion (Song et al. 2017), or more intuitively, completing and segmenting something unknown. Previous works on point cloud forecasting (Ilg et al. 2017; Dosovitskiy et al. 2015) often formulate it as a flow prediction problem where future frames are treated as the deformation of earlier partial observations. However, they cannot provide a complete scene understanding which is crucial for occlusion inference or robot decision making. On the other hand, works on semantic scene completion (Roldao, De Charette, and Verroust-Blondet 2022) consider only static scenes. Previous research treated forecasting and completion as two separate problems and no works exploited their connections deeply.
@@
Figure 2: The overview of SCSFNet. First, a 4D point cloud sequence will be transformed into 0.01m sparse voxels. The 4D sparse convolution encoder then generates spatial-temporal features at different scales (0.08m, 0.16m, and 0.32m). They are filled into 3D dense voxel grids and processed with attention-based skip connections at each scale. Finally, the model produces an implicit field based on a low-resolution voxel grid (0.08m), which is then used to generate a high-resolution voxel output (0.02m) filtered by the visibility grid for evaluation.
@@ 3.2 Attention-Based Skip Connection
At the most coarse-grained scale, dilated-convolution voxels are directly processed with up-convolutions to enter the next scale. At other scales, we combine dilated-convolution voxel features with positional embeddings, generating key-value pairs ${(K_{j},V_{j})}$ for key and value matrices $K$ and $V$ respectively. Positional embeddings are computed using Fourier basis of voxel centroid positions through an MLP. For the $i$-th up-convolution voxel at the previous scale, a query $Q_{i}$ is derived from its voxel feature and positional embedding. This allows us to compute the skipped feature for the $i$-th up-convolution voxel using $\text{softmax}(\frac{Q_{i}K^{T}}{\sqrt{d_{k}}})V$, where $d_{k}$ is the dimension of $Q_{i}$. In practice, we utilize multi-head cross-attentions and concatenate them to the final skipped features. Following U-Net’s common practice, we combine skipped features with original up-convolution voxel features, which are then forwarded to the decoder.
@@ 4.1 IGPLAY & IGNAV
While IGPLAY already exhibits many features suitable for 4D forecasting, the whole dynamics of the scene only comes from the viewer’s movement and interaction with the objects. To provide a new perspective of the underlying dynamics, we provide a new dataset called IGNAV with several active robots navigating in the scene. Each scene in IGNAV contains 9 semantic classes. We captured 600 scenes lasting 10 timesteps each, where several robots move around the scene in random speeds, with another robot watching them.
@@ A.2 Dataset Generation
To generate the simulated scenes in IGPLAY, we first specify the desired configuration of the environment in the logic language BDDL. In a typical BDDL description that we use, there is 1 floor, 2 tables, 1 bed, 2 sofas, 2 chairs, 3 floor lamps, 2 table lamps, and 1 iGibson robot. We then specify the desired logical state to make sure that everything is on the floor except table lamps which are on the table. After iGibson succeeds in creating the ideal scenario described by BDDL, we randomly discard some objects in the scene, and then add two types of interactive and movable objects at random locations in the scene: toys and pops. These objects will drop on the table, on the sofa, on the chair, or just on the ground. This process was repeated and finally, we captured 1000 scenarios, each lasting 10 timesteps, with various data captured from the robot which is interacting with toys among the furniture.
@@ B.2 Occlusion4D
In the training procedure, it samples solid points close to the target point cloud as positive instances and air points far away from the target point cloud as negative instances. Then it concatenates positive and negative instances to obtain query points, and trains the network to predict properties on these points including density, semantic label and so on based on the implicit representation of the scene.
@@ B.3 ST3DCNN
This baseline (Mersch et al. 2022) utilizes a novel 3D representation composed of concatenated range images as input. It simultaneously predicts future range images by a encoder-decoder structure and assigns per-point scores indicating validity or invalidity for multiple future time steps. We refer to this approach as ST3DCNN, which we consider as the SOTA approach (without sensor extrinsics) on the KITTIOdometry dataset, in line with the description from (Khurana et al. 2023).
@@ C Additional Ablation Studies
Supervising implicit functions through on-the-fly random samples is both effective and efficient. During the training phase, we employ random sampling of $10^{5}$ points from all the central coordinates of ground truth voxels. Another choice is using the entire ground truth, which entails calculating the loss based on $240\times 144\times 240=8294400$ voxels for a single data point in IGPLAY. This method, however, proves time-consuming and yields slightly inferior results (IoU drops from 56.5% to 56.3%). This substantiates that employing a random subset of $10^{5}$ voxels on-the-fly is significantly faster while maintaining effectiveness.
@@ Methods
In section 5.2 of the manuscript, we compared results between movable and static classes and report them in Table 2 in the manuscript. A class is labeled as movable if the corresponding objects can move, such as robots in IGNAV or cars in SemanticKITTI. We present a comprehensive breakdown of IoU values for each category and offer an in-depth analysis of the dynamics within each dataset.
@@ E Semantic Scene Completion tasks for SCSFNet
Our paper proposed the SCSF task, which mainly focuses on future forecasting. If we do not forecast and only focus on the current frame (the last given frame), degenerate cases for CSF (Complete Scene Forecasting) and SCSF (Semantic Complete Scene Forecasting) become SC (Scene Completion) and SSC (Semantic Scene Completion). And our SCSFNet can solve the SSC problem using the same pipeline for the SCSF task in different dimensions.

### 15 Dec 2023 v2

@@ 1 Introduction
The SCSF problem is very challenging since it requires both future forecasting and semantic scene completion (Song et al. 2017), or more intuitively, completing and segmenting something unknown. Previous works on point cloud forecasting (Ilg et al. 2017; Dosovitskiy et al. 2015) often formulate it as a flow prediction problem where future frames are treated as the deformation of earlier partial observations. However, they cannot provide a complete scene understanding which is crucial for occlusion inference or robot decision-making. On the other hand, works on semantic scene completion (Roldao, De Charette, and Verroust-Blondet 2022) consider only static scenes. Previous research treated forecasting and completion as two separate problems and no works exploited their connections deeply.
@@
Figure 2: The overview of SCSFNet. First, a 4D point cloud sequence $\mathcal{X}$ will be transformed into 0.01m sparse voxels. The 4D sparse convolution encoder then generates spatial-temporal features at different scales (0.08m, 0.16m, and 0.32m). They are filled into 3D dense voxel grids and processed with attention-based skip connections at each scale. Finally, the model produces an implicit field $\mathcal{F}$ based on a low-resolution voxel grid $E$ (0.08m), which is then used to generate a high-resolution voxel output (0.02m) filtered by the visibility grid for evaluation.
@@ 3.2 Attention-Based Skip Connection
At the most coarse-grained scale, dilated-convolution voxels are directly processed with up-convolutions to enter the next scale. At other scales, we combine dilated-convolution voxel features with positional embeddings, generating key-value pairs ${(K_{j},V_{j})}$ for key and value matrices $K$ and $V$ respectively. Positional embeddings are computed using the Fourier basis of voxel centroid positions through an MLP. For the $i$-th up-convolution voxel at the previous scale, a query $Q_{i}$ is derived from its voxel feature and positional embedding. This allows us to compute the skipped feature for the $i$-th up-convolution voxel using $\text{softmax}(\frac{Q_{i}K^{T}}{\sqrt{d_{k}}})V$, where $d_{k}$ is the dimension of $Q_{i}$. In practice, we utilize multi-head cross-attentions and concatenate them to the final skipped features. Following U-Net’s common practice, we combine skipped features with original up-convolution voxel features, which are then forwarded to the decoder.
@@ 4.1 IGPLAY & IGNAV
While IGPLAY already exhibits many features suitable for 4D forecasting, the whole dynamics of the scene only come from the viewer’s movement and interaction with the objects. To provide a new perspective of the underlying dynamics, we provide a new dataset called IGNAV with several active robots navigating in the scene. Each scene in IGNAV contains 9 semantic classes. We captured 600 scenes lasting 10 timesteps each, where several robots move around the scene in random speeds, with another robot watching them.
@@ A.2 Dataset Generation
To generate the simulated scenes in IGPLAY, we first specify the desired configuration of the environment in the logic language BDDL. In a typical BDDL description that we use, there is 1 floor, 2 tables, 1 bed, 2 sofas, 2 chairs, 3 floor lamps, 2 table lamps, and 1 iGibson robot. We then specify the desired logical state to make sure that everything is on the floor except table lamps which are on the table. After iGibson succeeds in creating the ideal scenario described by BDDL, we randomly discard some objects in the scene and then add two types of interactive and movable objects at random locations in the scene: toys and pops. These objects will drop on the table, on the sofa, on the chair, or just on the ground. This process was repeated and finally, we captured 1000 scenarios, each lasting 10 timesteps, with various data captured from the robot which is interacting with toys among the furniture.
@@ B.2 Occlusion4D
In the training procedure, it samples solid points close to the target point cloud as positive instances and air points far away from the target point cloud as negative instances. Then it concatenates positive and negative instances to obtain query points and trains the network to predict properties on these points including density, semantic label and so on based on the implicit representation of the scene.
@@ B.3 ST3DCNN
This baseline (Mersch et al. 2022) utilizes a novel 3D representation composed of concatenated range images as input. It simultaneously predicts future range images by an encoder-decoder structure and assigns per-point scores indicating validity or invalidity for multiple future time steps. We refer to this approach as ST3DCNN, which we consider as the SOTA approach (without sensor extrinsics) on the KITTIOdometry dataset, in line with the description from (Khurana et al. 2023).
@@ C Additional Ablation Studies
Supervising implicit functions through on-the-fly random samples is both effective and efficient. During the training phase, we employ random sampling of $10^{5}$ points from all the central coordinates of ground truth voxels. Another choice is using the entire ground truth, which entails calculating the loss based on $240\times 144\times 240=8294400$ voxels for a single data point in IGPLAY. This method, however, proves time-consuming and yields slightly inferior results (IoU drops from 56.5% to 56.3%). This substantiates that employing a random subset of $10^{5}$ voxels on the fly is significantly faster while maintaining effectiveness.
@@ Methods
In section 5.2 of the manuscript, we compared results between movable and static classes and reported them in Table 2 in the manuscript. A class is labeled as movable if the corresponding objects can move, such as robots in IGNAV or cars in SemanticKITTI. We present a comprehensive breakdown of IoU values for each category and offer an in-depth analysis of the dynamics within each dataset.
@@ E Semantic Scene Completion tasks for SCSFNet
Our paper proposed the SCSF task, which mainly focuses on future forecasting. If we do not forecast and only focus on the current frame (the last given frame), degenerate cases for CSF (Complete Scene Forecasting) and SCSF (Semantic Complete Scene Forecasting) become SC (Scene Completion) and SSC (Semantic Scene Completion). Our SCSFNet can solve the SSC problem using the same pipeline for the SCSF task in different dimensions.
@@ F Additional Visualizations
G Limitations and Future work
In our work, we require dynamic scenes that can support completion and forecasting at the same time. Given the limited options for real-world datasets meeting these criteria, we find SemanticKITTI to be the only suitable choice, which consists of large-scale 4D point cloud sequences with completion and semantic annotations. To further demonstrate the effectiveness of our method in the indoor setting, we generate 2 large-scale indoor datasets. We hope to extend the evaluation of our approach to suitable real-world datasets released in the future.
Additionally, in our experiments, we only performed relatively short-term forecasting due to the inherent challenges of the SCSF task. This limitation hinders the model’s ability to provide an accurate understanding of the entire scene, particularly focusing on dynamic objects, in the long term. In the future, we aspire to see more efficient approaches that can handle long-term SCSF tasks with minimal computational overhead.